Calculate the value of Kp for the equation

C(s)+CO2(g)−⇀↽−2CO(g)Kp=?

given that at a certain temperature

C(s)+2H2O(g) −⇀↽−CO2(g)+2H2(g)Kp1=3.73H2(g)+CO2(g) −⇀↽−H2O(g)+CO(g)Kp2=0.689
Answer #1

2 ( ) مع ( م مژه هات * ( و/ce 2 د Date Page. Givens Key = 3:22 C(s) +24019) = (0,19) +24, 6-0 Ep-clk59 HQ +CG,(9) = ,019) +6

Hope you like the solution.If you have any further queries,you can ask in the comment box and i will definitely have them solved.Please hit the like butoon,please.

> In the end instead of adding at the end, you need to multiply the value you got. I.E. 3.72x(0.689)^2. By multiplying instead of dividing you will obtain the correct answer. However, the rest of the process is correct.

Reece Rodriguez Tue, Mar 29, 2022 12:30 PM

> instead of adding excuse me.

Reece Rodriguez Tue, Mar 29, 2022 12:30 PM